The company wants to determine its re-order point(R). demand is variable and theh want to build a safety stock into R. if the average daily demands 16, desired time is 7 days, the desire to "z" value is 1.96 and a standard deviation of usage during lead time is 4, what is the desired value of R? round to the nearest integer
Answer #1

Solution:

The formula for Reorder point = Daily demand * L +(z*sd*\sqrt{L})

Daily demand = 16

Lead time, L = 7 days

z =1.96

sd = 4

R = (16*7) + (1.96 * 4 * \sqrt{7} )

R = 112 + ( 7.84 * 2.645)

R = 112 + 20.74

R = 132.74 = 133 units